# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin principle

### v2

@@ Abstract
We derive the law of entropy non-decrease directly from the Kelvin principle for simple and compound systems without using Clausius inequality. A key of the derivation is a formulation of entropy in terms of work by a Carnot machine operating between a system and a single heat reservoir, which is equivalent to Clausius entropy based on heat and Gyftopoulos-Beretta entropy based on work.
1 Introduction
How information of irreversibility is encoded into entropy as a state quantity? This would be one of the most difficult points for those who learn thermodynamics first to grasp. This is also true for physicists today, and the famous paper by Lieb and Yngvason was aimed to clarify the nature of entropy in terms of adiabatic accessibility in their axiomatic approach to thermodynamics [1], where neither statistical mechanics nor heat engines are used [2] (see also Ref. [3]).
@@ 1 Introduction
which is the law of entropy non-decrease, and the equality applies to reversible adiabatic processes. Here, the entropies of the system’s states $A_{1}$ and $A_{2}$ are measured by
@@
where $S_{0}$ is the entropy at a reference state $A_{0}$, $\delta Q$ denotes heat supplied to the system and $T_{R}$ denotes the temperature of a heat reservoir. During the process from $A_{0}$ to $A_{1}$, the heat reservoir that contacts the system and its temperature $T_{R}$ may continuously change. Equation (2) is the formula derived by Clausius. Usually, in standard textbooks on thermodynamics such as Ref. [4], after introducing Carnot cycle and Carnot efficiency as the maximum efficiency of heat engines, Eq. (1) is derived by using Clausius inequality for a cyclic process:
@@
The existence of entropy as a state quantity (2) is shown by using the equality condition of Clausius inequality. Then, by applying Clausius inequality Eq. (3) to a cyclic process $A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, where “rev” implies that the process is reversible, we have
$S_{2}-S_{1}-\int_{1}^{2}\frac{\delta Q}{T_{R}}\geq 0.$ (4)
If the process $A_{1}\to A_{2}$ is an adiabatic process, we may put $\delta Q=0$ into the above equality and derive Eq. (1).
The sketch of the above derivation of the law of entropy non-decrease is based on Clausius inequality Eq. (3). The Clausius inequality is derived from the Kelvin principle, which manifests one expression of the second law of thermodynamics: We cannot extract a positive work from a single heat reservoir at a fixed temperature without leaving any other changes. For a cyclic process, it can be expressed as
@@
where $W_{\rm cyc}$ denotes the work extracted from the cycle starting from an equilibrium state, and the equality applies to a reversible cycle. Deriving Clausius inequality from Kelvin principle relies on a complicated thermodynamic process tactically involving multiple heat reservoirs at different temperatures and multiple Carnot heat engines [4]. This is natural because the Kelvin principle originally refers to a process with a single heat reservoir. Moreover, the law of entropy non-decrease Eq. (1) is derived as a special case ($\delta Q=0$) of the Clausius inequality Eq. (3) and it is thus derived indirectly from the Kelvin principle.
@@
In this paper, we show that we can derive the law of entropy non-decrease Eq. (1) directly from Kelvin principle Eq. (5) without using Clausius inequality Eq. (3). To this end, we newly define the entropy of a substance in terms of the maximum work of a Carnot machine operating between the substance and a single heat reservoir. Because we will assume only a single heat reservoir in our derivation, it apparently makes it possible to apply the Kelvin principle to show the law of entropy non-decrease. We also discuss the relationship between our entropy and Clausius entropy (2) or Gyftopoulos-Beretta entropy [5] based on work.
2 Main results
@@
We derive the principle of entropy non-decrease Eq. (1) from Kelvin principle Eq. (5). After deriving our entropy in terms of work by a Carnot machine in Sec. 2.1, we give the derivation for simple systems in Sec 2.2 and compound systems in Sec. 2.3. Here, a simple system is a system in an equilibrium state that can be specified with the system’s temperature $T$ and volume $V$. A compound system is a system composed of simple systems.
2.1 Definition of entropy as state quantity
@@
Figure 1: A reversible path $\gamma$ from $A_{0}$ to $A_{1}$ and a reversible path $\gamma^{\prime}$ from $A_{1}$ to $A_{0}$ on the temperature-volume ($T$$V$) plane. They consist of the adiabatic processes (arrows along the adiabatic curves) and the temperature-changing processes with a Carnot machine operating between the system at constant volume and the heat reservoir at $T_{0}$ (vertical arrows), respectively.
@@
We introduce entropy as a state quantity in our framework. Consider a simple system initially in an equilibrium state $A_{0}$ specified with $(T_{0},V_{0})$ with $T_{0}$ and $V_{0}$ the system’s initial temperature and volume, respectively. Consider a reversible path $\gamma$ from the equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$ and a reversible path $\gamma^{\prime}$ from the state $A_{1}$ to $A_{0}$ (Fig. 1). Any reversible path $\gamma$ from the equilibrium state $A_{0}$ to another one $A_{1}$ can be realized by a combination of (i) an adiabatic process (the arrow along the adiabatic curve in Fig. 1) and (ii) a temperature-changing process with a Carnot machine operating between the system at constant volume and the single heat reservoir at $T_{0}$ to raise or lower the system’s temperature (the vertical arrow in Fig. 1) by regarding the system as a finite-sized heat reservoir; any reversible path $\gamma$ may be decomposed into $N$-reversible processes $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$ with $\gamma_{i}$ being the $i$-th reversible process, which is either of the above two processes (i) and (ii). In particular, when the path is a continuous one, it may be realized by a combination of the thermodynamic processes (i) and (ii) by taking the limit of $N\to\infty$.
@@
$W_{0\to 1}^{\rm rev}[\gamma]=U_{0}-U_{1}+T_{0}\int_{\gamma}\frac{\delta Q}{T},$ (6)
$W_{1\to 0}^{\rm rev}[\gamma^{\prime}]=U_{1}-U_{0}+T_{0}\int_{\gamma^{\prime}}\frac{\delta Q}{T},$ (7)
where $U$ is the internal energy and $T$ is the temperature of the system, not the temperature of the heat reservoir; the heat reservoir at fixed temperature $T_{R}=T_{0}$ is used here. We will derive Eqs. (6) and (7) later.
By applying the equality case of Kelvin principle Eq. (5) to the reversible cycle $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
@@
where $-\gamma^{\prime}$ denotes the reversed path of $\gamma^{\prime}$. Because $\gamma$ and $-\gamma^{\prime}$ are arbitrary paths connecting $A_{0}$ and $A_{1}$, Eq. (9) imply that the quantity $\int_{\gamma}\delta Q/T$ has a unique value independent of a path, which we denote by $\int_{0}^{1}\delta Q/T$. Therefore, we can define a state quantity $S_{1}$ at the state $A_{1}$ as
@@
with $S_{0}$ at the state $A_{0}$ being a criterion value. This state quantity $S$ is our entropy. We can write the reversible work $W_{0\to 1}^{\rm rev}$ from $A_{0}$ to $A_{1}$ in Eq. (6) as
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{0}(S_{0}-S_{1})$ (11)
in a form independent of a path $\gamma$ by using the entropy. This quantity is often referred to as exergy as the maximum work extracted from a finite-sized heat reservoir until it reaches an equilibrium state with a heat reservoir at fixed temperature [6]. It should be noted that the temperature $T$ in Eq. (10) is the system’s temperature, not the heat reservoir’s one as in Eq. (2). While Eq. (2) is usually obtained via the equality case of Clausius inequality assuming thermal contact with infinitely many heat reservoirs with different temperatures, Eq. (10) has been obtained by directly applying Kelvin principle assuming only a single heat reservoir at fixed temperature $T_{0}$. But, once Eq. (10) is established, we may interpret that it is equivalent to the Clausius entropy Eq. (2), as any path in Fig. 1 can be realized with a reversible process during which a system contacts infinitely many heat reservoirs with different temperatures, and the system’s temperature $T$ agrees with the temperature $T_{R}$ of each heat reservoir for such a reversible process.
Now, we derive Eqs. (6) and (7). The work $\Delta W_{i}$ extracted from the $i$-th reversible process $\gamma_{i}$ from $(T^{(i)},V^{(i)})$ to $(T^{(i+1)},V^{(i+1)})$ as the part of $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$, where $T^{(0)}=T_{0}$ and $T^{(N)}=T_{1}$, is given by the following two cases:
@@
$({\rm ii})\ \Delta W_{i}^{\rm rev}[\gamma_{i}]$ $=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T},$ (13)
depending on the type of the $i$-th process, where $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is independent of $\gamma_{i}$ and $V^{(i+1)}=V^{(i)}$ (fixed volume) in Eq. (13) for the process (ii). These works (12) and (13) are obtained as follows: The work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is just given by the internal energy change $U^{(i)}-U^{(i+1)}=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})$. For the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) for the process (ii), we operate a Carnot machine (heat engine, refrigerator, or heat pump) between the system at temperature $T$ and the heat reservoir at temperature $T_{0}$ until the system’s temperature $T$ initiating from $T^{(i)}$ reaches $T^{(i+1)}$ by keeping $V=V^{(i)}$. To raise the system’s temperature $T$, we may operate a Carnot heat engine for $T<T_{0}$ and a Carnot heat pump for $T_{0}\leq T$; to lower the system’s temperature $T$, we may operate a Carnot heat engine for $T_{0}<T$ and a Carnot refrigerator for $T\leq T_{0}$. For any case, the infinitesimal work $\delta W$ extracted by a Carnot machine operating between the system at $T$ and the heat reservoir at $T_{0}$ is commonly given by
$\delta W=-\left(1-\frac{T_{0}}{T}\right)\delta Q,$ (14)
@@
$\Delta W_{i}^{\rm rev}[\gamma_{i}]=\int_{\gamma_{i}}\delta W$ $=-\int_{\gamma_{i}}\left(1-\frac{T_{0}}{T}\right)\delta Q$ (15)
$=-\int_{\gamma_{i}}\delta Q+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T}$
$=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\gamma_{i}}\frac{\delta Q}{T}.$
@@
$=\sum_{i=0}^{N-1}U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{0}\int_{\sum_{i=0}^{N-1}\gamma_{i}}\frac{\delta Q}{T}$
$=U_{0}-U_{1}+T_{0}\int_{\gamma}\frac{\delta Q}{T},$
@@
2.2 Simple systems
Figure 2: An arbitrary adiabatic process from an equilibrium state $A_{1}$ of a simple system to another equilibrium state $A_{2}$ (dashed curve), where $A_{1}$ and $A_{2}$ are created reversibley from a reference equilibrium state $A_{0}$ (solid curves). The Kelvin principle Eq. (5) is applied to the cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$.
Now that we have derived the entropy in Eq. (10), we are ready to show its non-decrease during adiabatic processes. That is, we show Eq. (1) from Eq. (5). Here, we show it for simple systems.
Let a thermally isolated system be in the equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$. Consider an arbitrary adiabatic process from $A_{1}$ to another equilibrium state $A_{2}$ with $(T_{2},V_{2})$ (Fig. 2) by adiabatic operations such as stirring and moving a piston attached to the system, during which the system’s state may be in nonequilibrium; after the operations, we may wait for a sufficiently long time until the system reaches the equilibrium state $A_{2}$. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by

### v5

# Formulation of entropy through work by Carnot machine and direct derivation of law of entropy non-decrease from Kelvin-Planck principle

@@ Abstract
We derive the law of entropy non-decrease directly from the Kelvin-Planck principle for simple and compound systems without using the Clausius inequality. A key of the derivation is a formulation of entropy in terms of work by a Carnot machine operating between a system and a single heat reservoir at fixed temperature, which is equivalent to Clausius entropy based on heat and Gyftopoulos-Beretta entropy based on work. We also show that we can characterize entropy as the thermodynamic cost to create a system’s state from a reference state.
I Introduction
How information of irreversibility is encoded into entropy as a state quantity? This would be one of the most difficult points for those who learn thermodynamics first to grasp. This is also true for physicists today, and the famous paper by Lieb and Yngvason was aimed to elucidate the nature of entropy in terms of adiabatic accessibility in their axiomatic approach to thermodynamics [1, 2, 3], where neither statistical mechanics nor heat engines are used [3] (see also Ref. [4]).
@@ I Introduction
which is the law of entropy non-decrease, and the equality applies to reversible adiabatic processes. Here, the entropies of the system’s states $A_{1}$ and $A_{2}$ are given by
@@
where $S_{0}$ is the entropy at a reference state $A_{0}$, $\delta Q$ denotes heat supplied to the system, and $T_{R}$ denotes the temperature of a heat reservoir. During the process from $A_{0}$ to $A_{1}$, the temperature $T_{R}$ of the heat reservoir that contacts the system may change [5]. Equation (2) is the formula derived by Clausius. Usually, in standard textbooks on thermodynamics such as Ref. [6], after introducing Carnot cycle and Carnot efficiency as the maximum efficiency of heat engines, Eq. (1) is derived by using the Clausius inequality for a cyclic process:
@@
The existence of entropy as a state quantity (2) is shown by using the equality condition of the Clausius inequality. Then, by applying the Clausius inequality Eq. (3) to a cyclic process $A_{1}\to A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}$, where “rev” implies that the process is reversible, we have
$S_{2}-S_{1}-\int_{\Gamma}\frac{\delta Q}{T_{R}}\geq 0.$ (4)
Here, $\Gamma$ denotes a path from $A_{1}$ to $A_{2}$. If the process $A_{1}\to A_{2}$ is an adiabatic process, which we write as $A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}$, we may put $\delta Q=0$ into the above equality and derive Eq. (1).
The sketch of the above derivation of the law of entropy non-decrease is based on the Clausius inequality Eq. (3). The Clausius inequality is derived from the Kelvin-Planck principle, which manifests one expression of the second law of thermodynamics: We cannot extract a positive work from a cyclic process with a single heat reservoir at fixed temperature:
@@
where $W_{\rm cyc}$ denotes the work extracted from a cycle starting from an equilibrium state, and the equality applies to a reversible cycle. Deriving the Clausius inequality from the Kelvin-Planck principle relies on a complicated thermodynamic process that tactically involves multiple (or infinitely many) heat reservoirs at different temperatures and multiple (or infinitely many) Carnot heat engines [6]. This is because the Kelvin-Planck principle is originally applied to a process with a single heat reservoir at fixed temperature, and a complicated setup is necessary for it to be applied to a case with multiple (or infinitely many) heat reservoirs. Moreover, the law of entropy non-decrease Eq. (1) is derived as a special case of the Clausius inequality Eq. (3) (i.e., the case of $\delta Q=0$ in Eq. (4)) and it is thus derived indirectly from the Kelvin-Planck principle.
@@
In this paper, we show that we can derive the law of entropy non-decrease Eq. (1) directly from the Kelvin-Planck principle Eq. (5) without using the Clausius inequality Eq. (3). To this end, we give a new formulation of entropy of a substance in terms of the work of a Carnot machine operating between the substance and a single heat reservoir at fixed temperature. Because we will assume only a single heat reservoir at fixed temperature in our derivation, it apparently makes it possible to apply the Kelvin-Planck principle to show the law of entropy non-decrease. We discuss the relationship between our entropy and the Clausius entropy (2) based on heat or the Gyftopoulos-Beretta entropy [7] based on work. Moreover, we also show that entropy may be characterized by the thermodynamic cost to create a system’s state from a reference state.
II Main results
@@
We derive the law of entropy non-decrease Eq. (1) from the Kelvin-Planck principle Eq. (5). After the formulation of our entropy in terms of work by a Carnot machine in Sec. II.1, we give the derivation for simple systems in Sec II.2 and compound systems in Sec. II.3. Here, a simple system is a system in an equilibrium state that can be specified with the system’s temperature $T$ and volume $V$. A compound system is a system composed of simple systems.
II.1 Formulation of entropy as state quantity
@@
Figure 1: A reversible path $\gamma$ from $A_{0}$ to $A_{1}$ and a reversible path $\gamma^{\prime}$ from $A_{1}$ to $A_{0}$ on the temperature-volume ($T$$V$) plane. They consist of the reversible adiabatic processes (arrows along the adiabatic curves) and the reversible temperature-changing processes with a Carnot machine operating between a system at constant volume and a heat reservoir at temperature $T_{0}$ (vertical arrows).
@@
We introduce our entropy as a state quantity in the following way. Consider a simple system initially in an equilibrium state $A_{0}$ specified with $(T_{0},V_{0})$ with $T_{0}$ and $V_{0}$ the system’s initial temperature and volume, respectively. Consider a reversible path $\gamma$ from the equilibrium state $A_{0}$ to another equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$ and a reversible path $\gamma^{\prime}$ from the state $A_{1}$ to $A_{0}$ (Fig. 1). Any reversible path $\gamma$ from the equilibrium state $A_{0}$ to another one $A_{1}$ can be realized by a combination of (i) a reversible adiabatic process (the arrow along the adiabatic curve in Fig. 1) and (ii) a reversible temperature-changing process with a Carnot machine operating between the system at constant volume and a single heat reservoir at fixed temperature $T_{o}$ to raise or lower the system’s temperature (the vertical arrow in Fig. 1) by regarding the system as a finite-sized heat reservoir [5]; any reversible path $\gamma$ may be decomposed into $N$-reversible processes $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$ with $\gamma_{i}$ being the $i$-th reversible process, which is either of the above two thermodynamic processes (i) and (ii). In particular, when the path is continuous, it may be realized by a combination of the processes (i) and (ii) by taking the limit of $N\to\infty$. It should be noted that there is no need for the system’s temperature to match the fixed temperature $T_{o}$ of the heat reservoir during the process including the initial and final states $A_{0}$ and $A_{1}$.
@@
$W_{0\to 1}^{\rm rev}[\gamma]=U_{0}-U_{1}+T_{o}\int_{\gamma}\frac{\delta Q}{T},$ (6)
and
$W_{1\to 0}^{\rm rev}[\gamma^{\prime}]=U_{1}-U_{0}+T_{o}\int_{\gamma^{\prime}}\frac{\delta Q}{T},$ (7)
where $U$ is the internal energy of the system and $T$ is the temperature of the system, not the temperature of the heat reservoir; the heat reservoir at fixed temperature $T_{o}$ is used here. We will derive Eqs. (6) and (7) later.
By applying the equality case of the Kelvin-Planck principle Eq. (5) to the reversible cycle $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$, we have
@@
where $-\gamma^{\prime}$ denotes the reversed path of $\gamma^{\prime}$. Because $\gamma$ and $-\gamma^{\prime}$ are arbitrary paths connecting $A_{0}$ and $A_{1}$, Eq. (9) implies that the quantity $\int_{\gamma}\delta Q/T$ has a unique value independent of a path, which we denote by $\int_{0}^{1}\delta Q/T$. Therefore, we can define a state quantity $S$, which takes $S_{1}$ at the state $A_{1}$ as
@@
with $S_{0}$ at the state $A_{0}$ a criterion value. This state quantity $S$ is our entropy. We can write the reversible work $W_{0\to 1}^{\rm rev}$ from $A_{0}$ to $A_{1}$ in Eq. (6) as
$W_{0\to 1}^{\rm rev}=U_{0}-U_{1}-T_{o}(S_{0}-S_{1})$ (11)
in a form independent of a path $\gamma$ by using the entropy. Equation (11) serves as the maximum work extracted from a system with the aid of a heat reservoir at fixed temperature (see Appendix A for a simple proof based on Ref. [8]). It should be noted again that the temperature $T$ in Eq. (10) is the system’s temperature, not the heat reservoir’s one as in Eq. (2). While Eq. (2) is usually obtained via the equality case of the Clausius inequality Eq. (3) that assumes thermal contact with multiple (or infinitely many) heat reservoirs with different temperatures, Eq. (10) has been obtained by directly applying the Kelvin-Planck principle assuming only a single heat reservoir at fixed temperature $T_{o}$. But, once Eq. (10) is established, we may interpret that it is equivalent to the Clausius entropy Eq. (2), as any path in Fig. 1 can be realized with a reversible process during which a system contacts multiple (or infinitely many) heat reservoirs with different temperatures, and the system’s temperature $T$ agrees with the temperature $T_{R}$ of each heat reservoir for such a reversible process.
Now, we derive Eqs. (6) and (7). The reversible work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ extracted from the $i$-th reversible process $\gamma_{i}$ from $(T^{(i)},V^{(i)})$ to $(T^{(i+1)},V^{(i+1)})$ as the part of $\gamma=\sum_{i=0}^{N-1}\gamma_{i}$, where $T^{(0)}=T_{0}$ and $T^{(N)}=T_{1}$, is given by the following two cases:
@@
$({\rm ii})\ \Delta W_{i}^{\rm rev}[\gamma_{i}]$ $=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{o}\int_{\gamma_{i}}\frac{\delta Q}{T},$ (13)
depending on the type of the $i$-th process. Here, $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is independent of $\gamma_{i}$ (i.e., determined only by the initial and final states), and note that $U(T^{(i+1)},V^{(i+1)})=U(T^{(i+1)},V^{(i)})$ in Eq. (13) because of $V^{(i+1)}=V^{(i)}$ (fixed volume) for the process (ii). These works (12) and (13) are obtained as follows: The work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (12) for the process (i) is just given by the internal energy change $U^{(i)}-U^{(i+1)}=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})$. For the work $\Delta W_{i}^{\rm rev}[\gamma_{i}]$ in Eq. (13) for the process (ii), we operate a Carnot machine (heat engine, refrigerator, or heat pump) between the system at temperature $T$ and the heat reservoir at temperature $T_{o}$ until the system’s temperature $T$ initiating from $T^{(i)}$ reaches $T^{(i+1)}$ by keeping $V=V^{(i)}$. To raise the system’s temperature $T$, we may operate a Carnot heat engine for $T<T_{o}$ and a Carnot heat pump for $T_{o}\leq T$; to lower the system’s temperature $T$, we may operate a Carnot heat engine for $T_{o}<T$ and a Carnot refrigerator for $T\leq T_{o}$. For any case, the infinitesimal work $\delta W$ extracted by a Carnot machine operating between the system at $T$ and the heat reservoir at $T_{o}$ is commonly given by
$\delta W=-\left(1-\frac{T_{o}}{T}\right)\delta Q,$ (14)
@@
$\Delta W_{i}^{\rm rev}[\gamma_{i}]=\int_{\gamma_{i}}\delta W$ $=-\int_{\gamma_{i}}\left(1-\frac{T_{o}}{T}\right)\delta Q$ (15)
$=-\int_{\gamma_{i}}\delta Q+T_{o}\int_{\gamma_{i}}\frac{\delta Q}{T}$
$=U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{o}\int_{\gamma_{i}}\frac{\delta Q}{T}.$
@@
$=\sum_{i=0}^{N-1}U(T^{(i)},V^{(i)})-U(T^{(i+1)},V^{(i+1)})+T_{o}\int_{\sum_{i=0}^{N-1}\gamma_{i}}\frac{\delta Q}{T}$
$=U_{0}-U_{1}+T_{o}\int_{\gamma}\frac{\delta Q}{T},$
@@
II.2 Derivation for simple systems
Figure 2: An arbitrary adiabatic process from one equilibrium state $A_{1}$ of a simple system to another equilibrium state $A_{2}$ (dashed curve), where $A_{1}$ and $A_{2}$ are created reversibley from a reference equilibrium state $A_{0}$ (solid curves). The Kelvin-Planck principle Eq. (5) is applied to the cyclic process $A_{0}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{1}\stackrel{{\scriptstyle\rm ad}}{{\longrightarrow}}A_{2}\stackrel{{\scriptstyle\rm rev}}{{\longrightarrow}}A_{0}$.
Now that we have formulated our entropy in Eq. (10), we are ready to show its non-decrease during adiabatic processes. That is, we show Eq. (1) from Eq. (5). Here, we show it for simple systems.
Let a thermally isolated system be in the equilibrium state $A_{1}$ specified with $(T_{1},V_{1})$. Consider an arbitrary adiabatic process from $A_{1}$ to another equilibrium state $A_{2}$ with $(T_{2},V_{2})$ (Fig. 2) by adiabatic operations such as stirring and moving a piston attached to the system, during which the system may be in a nonequilibrium state; after the operations, we may wait for a sufficiently long time until the system reaches the equilibrium state $A_{2}$. The work extracted from this process $W_{1\to 2}^{\rm ad}$ is given by